# Correspondence-Free, Function-Based Sim-to-Real Learning for Deformable Surface Control

### v2

@@ I-C Our Method
The focus of this paper is a new learning-based sim-to-real transfer method that corrects the shape errors produced by NN-based forward prediction trained on simulation data, thereby enabling efficient and accurate NN-based IK computation for deformation control of soft robots, rather than proposing a new control strategy.
@@ II-C NN-based Deformation Function Space
$\color[rgb]{0,0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.0}\mathbf{p}_{i}^{*}=\Phi_{\bm{\gamma}}(\mathbf{p})=\Phi_{\mathcal{N}_{rbf}(\mathcal{S}_{j}^{c})}\left(\mathbf{B}(u_{i},v_{i},\mathcal{S}_{j}^{c})\right).$}} (4)
@@ V-A4 Compressible Soft Manipulator
The four hardware setups were selected to span representative deformation modes and structural complexities in pneumatically actuated soft robots. The deformable membrane provides inflation-induced deformations with regular chamber geometry, while the robotic mannequin involves inflation through freeform chamber boundaries with overlaps, resulting in more complex and spatially coupled surface deformations. The two soft manipulators highlight structural bending, with one driven purely by bending and the other combining inflation / compression and bending, thereby introducing hybrid deformation behaviors. Together, these systems cover the spectrum from simple to complex chamber inflation and from pure bending to coupled bending, inflation and compression, demonstrating that our correspondence-free sim-to-real learning framework generalizes across diverse actuation principles and deformation characteristics rather than being tailored to a single morphology.
@@ V-C Results of Sim-to-Real Learning by 3D Scanner
$\color[rgb]{0,0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.0}\mathcal{L}_{direct}=D_{\text{Haus}}+5.0D_{\text{Cham}}.$}} (20)

### v3

@@ I-C Our Method
This paper focuses on a learning-based sim-to-real transfer method that corrects shape errors in NN-based forward prediction trained on simulation data. The method establishes a function-based mapping between simulated and real deformation responses, enabling efficient and accurate NN-based IK computation for soft robot control. It builds on a learning-based theoretical foundation rather than a new control strategy.
@@ II-C NN-based Deformation Function Space
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbf{p}_{i}^{*}=\Phi_{\bm{\gamma}}(\mathbf{p})=\Phi_{\mathcal{N}_{rbf}(\mathcal{S}_{j}^{c})}\left(\mathbf{B}(u_{i},v_{i},\mathcal{S}_{j}^{c})\right).$}} (4)
@@ V-A4 Compressible Soft Manipulator
The four hardware setups were selected to span representative deformation modes and structural complexities in pneumatically actuated soft robots. The deformable membrane provides inflation-induced deformations with regular chamber geometry, while the robotic mannequin involves inflation through freeform chamber boundaries with overlaps, resulting in more complex and spatially coupled surface deformations. The two soft manipulators highlight structural bending, with one driven purely by bending and the other combining inflation / compression and bending, thereby introducing hybrid deformation behaviors. Together, these systems cover the spectrum from simple to complex chamber inflation and from pure bending to coupled bending, inflation and compression, demonstrating that our sim-to-real learning framework generalizes across diverse actuation principles and deformation characteristics rather than being tailored to a single morphology.
@@ V-C Results of Sim-to-Real Learning by 3D Scanner
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{L}_{direct}=D_{\text{Haus}}+5.0D_{\text{Cham}}.$}} (20)
@@ A-B Additional IK Tests with More Target Shapes
[Uncaptioned image] Yingjun Tian received the B.E. degree in mechanical engineering from the University of Science and Technology of China, Hefei, China, in 2019, and the Ph.D. degree in mechanical engineering from The University of Manchester, Manchester, U.K., in 2024. He is currently a Postdoctoral Research Associate with the Digital Manufacturing Lab, The University of Manchester. His research interests include computational design, soft robotics, and robot learning.
[Uncaptioned image] Guoxin Fang (Member, IEEE) received the B.E. degree in mechanical engineering from the Beijing Institute of Technology, Beijing, China, in 2016, and the Ph.D. degree in advanced manufacturing from Delft University of Technology, Delft, The Netherlands, in 2022. He is currently an Assistant Professor with the Department of Mechanical and Automation Engineering at The Chinese University of Hong Kong, Hong Kong. Prior to this, he was a Research Associate with the Department of Mechanical, Aerospace and Civil Engineering at The University of Manchester, U.K. His research interests include geometric computing, computational design, digital fabrication, and robotics.
[Uncaptioned image] Renbo Su received his B.Eng. degree in Material Forming and Control Engineering from East China University of Science and Technology, Shanghai, China, in 2019. He then earned his M.Sc. degree in Advanced Manufacturing Technology and Systems Management in 2020 and his Ph.D. degree in Mechanical Engineering in 2025, both from the University of Manchester, UK. He is currently a postdoctoral research fellow with the College of Intelligent Robotics and Advanced Manufacturing, Fudan University, Shanghai, China. His research interests include the design and manufacturing of rigid–soft hybrid robots, metamaterials, as well as the development of equipment for multi-axis additive manufacturing.
[Uncaptioned image] Aoran Lyu is currently pursuing a Ph.D. degree in Production and Manufacturing Engineering in the Department of Mechanical and Aerospace Engineering at the University of Manchester. He received his B.S. degree in Mathematics from South China University of Technology in 2021 and his M.E. degree in Computer Technology from the same university in 2024. His research interests include computer graphics, physics-based simulation, and computational design.
[Uncaptioned image] Neelotpal Dutta received the B.Tech. (Hons.) degree in Mechanical Engineering from the Indian Institute of Technology Mandi, India, in 2020, where he was awarded the President of India Gold Medal. He is currently a Ph.D. student in the Department of Mechanical and Aerospace Engineering at the University of Manchester, U.K. His research interests include geometric computing, digital manufacturing, and computer-aided design and manufacturing, with a focus on toolpath generation and process planning for additive and subtractive manufacturing.
[Uncaptioned image] Weiming Wang received the BS and PhD degrees from the Dalian University of Technology in 2010 and 2016, respectively. He was a Postdoctoral Researcher supported by the Marie Skłodowska-Curie LEaDing Fellows Programme at the Delft University of Technology, Delft, The Netherlands. Currently, he is a Postdoctoral Research Associate at The University of Manchester, Manchester, U.K. His research interests are Computational Fabrication and Additive Manufacturing.
[Uncaptioned image] Simeon Gill received his PhD on functional ease in 2009 at Manchester Metropolitan University (MMU) and has since focused on developing pattern-cutting theory informed by body scan data. He is currently a Reader in Fashion Technology at the University of Manchester, and his recent work uses parametric pattern techniques to automate engineered patterns that meet the fit and functional requirements of individual bodies.
[Uncaptioned image] Andrew Weightman graduated in 2006 with a Ph.D. in mechanical engineering from the University of Leeds. While at the University of Leeds, he developed rehabilitation robotic technology for improving upper limb function in adults and children with neurological impairment, which was successfully utilized in homes, schools, and clinical settings. He has been at The University of Manchester as part of the centre for robotics and artificial intelligence since 2013. He has research interests in biomimetic mobile robotics, rehabilitation robotics, robotics for nuclear decommissioning, and soft robotics.
[Uncaptioned image] Charlie C.L. Wang is currently Professor of Smart Manufacturing at the University of Manchester (UoM). Before joining UoM in 2020, he worked as Professor and Chair of Advanced Manufacturing at Delft University of Technology, The Netherlands and as Professor of Mechanical and Automation Engineering at the Chinese University of Hong Kong. He received his Ph.D. degree (2002) in mechanical engineering from Hong Kong University of Science and Technology (HKUST). His research interests include Digital Manufacturing, Computational Design, Additive Manufacturing, Soft Robotics, Geometric Computing, and Computer Graphics. He is Fellow of the American Society of Mechanical Engineers (ASME) and the Solid Modelling Association (SMA).